Page 1
CBSE BOARD
SAMPLE
PAPER
2026
SOLUTIONS
DOWNLOAD
Page 2
MATHEMATICS – Code No. 041
MARKING SCHEME
CLASS – XII (2025-26)
SECTION-A (MCQs of 1 mark each)
Sol. N. Hint / Solution Marks
1 1
Clearly from the graph Domain is [− 2 , 2]
1 So graph is of the function sin−1(2𝑥) 1
Answer is (B) sin−1(2𝑥)
1 1
Domain is [− 3 , 3]
1
So the function is cos −1 (3𝑥) 1
(V.I.)
Answer is (C) cos−1 (3𝑥)
AB is defined so n=4
AC is defined so p=4
2 AB and AC are square matrices of same order 1
so 𝑚 × 3 = 𝑚 × 𝑞 ⇒ 𝑞 = 3 = 𝑚
Answer is (A) 𝑚 = 𝑞 = 3 𝑎𝑛𝑑 𝑛 = 𝑝 = 4
As A is skew symmetric
So 𝑝 = 0, 𝑞 = 2, 𝑟 = −3, 𝑡 = 4
𝑞+𝑡 6
3 So 𝑝+𝑟 = −3 = −2 1
Answer is (A) -2
|𝑎𝑑𝑗 𝐴| = 27 ⇒ |𝐴|3 = 27 = 33 ⇒ |𝐴| = 3
4 𝐴 (𝑎𝑑𝑗 𝐴) = |𝐴| 𝐼 = 3 𝐼 1
Answer is (C) 𝟑 𝑰
1
0 0
3 0 0 3
1
Inverse of the matrix [0 2 0] = 0 2 0
5 0 0 5 1 1
[0 0 5]
Answer is (B)
cos 67𝑜 sin 67𝑜
| | = cos 67𝑜 cos 23𝑜 − sin 67𝑜 sin 23𝑜 = cos(67𝑜 + 23𝑜 ) =
sin 23𝑜 cos 23𝑜
6 cos 90𝑜 = 0 1
Answer is (A) 0
𝑓(𝑥) is continuous at 𝑥 = 𝜋
⇒ lim−(𝑘𝑥 + 1) = lim+ 𝑐𝑜𝑠 𝑥 = 𝑓(𝜋)
𝑥→𝜋 𝑥→𝜋
⇒ lim [𝑘(𝜋 − ℎ) + 1] = lim 𝑐𝑜𝑠( 𝜋 + ℎ ) = 𝑘𝜋 + 1
7 ℎ→0 ℎ→0 1
−2
⇒ 𝑘𝜋 + 1 = −1 ⇒𝑘 = 𝜋
−𝟐
Answer is (D) 𝝅
Page 1 of 12
Page 3
𝑓(𝑥) = 𝑥 tan−1 𝑥
1
𝑓 ′ (𝑥) = 1. tan−1 𝑥 + 𝑥 . 1+𝑥 2
8 1 𝜋 1 1
𝑓 ′ (1) = 1 . tan−1 1 + 1+1 = 4 +2
𝝅 𝟏
Answer is (B) 𝟒 + 𝟐
𝑓(𝑥) = 10 − 𝑥 − 2𝑥 2
⇒ 𝑓 ′ (𝑥) = −1 − 4𝑥
For increasing function 𝑓 ′ (𝑥) ≥ 0
⇒ − ( 1 + 4𝑥) ≥ 0
9 ⇒ ( 1 + 4𝑥) ≤ 0 1
⇒ 𝑥 ≤ − 1⁄4
1
⇒ 𝑥 ∈ (−∞, − 4]
𝟏
Answer is (A) (−∞, − ]
𝟒
𝑥𝑑𝑥 + 𝑦𝑑𝑦 = 0
⇒ ∫ 𝑥𝑑𝑥 = − ∫ 𝑦𝑑𝑦
𝑥2 𝑦2
⇒ 2 =− 2 +𝑘
10 1
⇒ 𝑥 2 + 𝑦 2 = 2𝑘
Solution is 𝑥 2 + 𝑦 2 = 2𝑘, 𝑘 being an arbitrary constant.
Answer is (C) Circles
𝑏 𝑏
𝐼 = ∫𝑎 𝑥 𝑓(𝑥)𝑑𝑥 = ∫𝑎 (𝑎 + 𝑏 − 𝑥)𝑓(𝑎 + 𝑏 − 𝑥)𝑑𝑥
𝑏
⇒ 𝐼 = ∫𝑎 (𝑎 + 𝑏 − 𝑥)𝑓(𝑥)𝑑𝑥 (given 𝑓(𝑎 + 𝑏 − 𝑥) = 𝑓(𝑥) )
𝑏 𝑏
⇒ 𝐼 = ∫𝑎 (𝑎 + 𝑏) 𝑓(𝑥)𝑑𝑥 − ∫𝑎 𝑥 𝑓(𝑥)𝑑𝑥
11 𝑏
⇒ 2 𝐼 = (𝑎 + 𝑏) ∫𝑎 𝑓(𝑥)𝑑𝑥 1
1 𝑏
⇒ 𝐼 = 2 (𝑎 + 𝑏) ∫𝑎 𝑓(𝑥)𝑑𝑥
𝒂+𝒃 𝒃
Answer is (D) 𝟐 ∫𝒂 𝒇(𝒙)𝒅𝒙
Let 𝐼 = ∫ 𝑥 3 𝑠𝑖𝑛4 (𝑥 4 ) cos(𝑥 4 ) 𝑑𝑥
1
Let sin(𝑥 4 ) = 𝑡 ⇒ 4𝑥 3 cos(𝑥 4 ) 𝑑𝑥 = 𝑑𝑡 ⇒ 𝑥 3 cos(𝑥 4 ) = 4 𝑑𝑡
1 1 1
Thus 𝐼 = ∫ 𝑡 4 ( 4 𝑑𝑡 ) = 20 𝑡 5 + C = 20 𝑠𝑖𝑛5 (𝑥 4 ) + C
12 1 1
⇒ 𝐼 = 20 𝑠𝑖𝑛5 (𝑥 4 ) + C = 𝑎 𝑠𝑖𝑛5 (𝑥 4 ) + C
1
So, 𝑎 = 20
𝟏
Answer is (B) 𝟐𝟎
The projection of the vector 𝑖̂ + 2𝑗̂ + 𝑘̂ on the line
1x1 +2x2+1x3 8
𝑟⃗ = (3𝑖̂ − 𝑗̂) + λ(𝑖̂ + 2𝑗̂ + 3𝑘̂ ) is = units
13 √12 2 2 +2 +3 √14 1
𝟖
Answer is (C) units
√𝟏𝟒
Page 2 of 12
Page 4
The distance of the point (a, b, c) from the y-axis is √𝑎2 + 𝑐 2
14 So, the distance is √32 + 52 = √34 units. 1
Answer is (B) √𝟑𝟒 units
(2𝑎⃗. 𝑖̂)𝑖̂ − (𝑏⃗⃗. 𝑗̂)𝑗̂ + (𝑐⃗. 𝑘̂)𝑘̂ = (2 × 3)𝑖̂ − (1)𝑗̂ + (2)𝑘̂
15 = 6𝑖̂ − 𝑗̂ + 2𝑘̂ = 𝑐⃗ 1
Answer is (D) 𝒄
⃗⃗
The points (1,0) and (0,2) satisfy the equation 2𝑥 + 𝑦 = 2
And shaded region shows that (0,0) doesn’t lie in the feasible solution region
16 So, the inequality is 2𝑥 + 𝑦 ≥ 2 1
Answer is (B) 𝟐𝒙 + 𝒚 ≥ 𝟐
(4,0) and (0,3) gives maximum value so
16 𝑍(4,0) = 𝑍(0,3) ⇒ 4𝑎 + 𝑐 = 3𝑏 + 𝑐 ⇒ 4𝑎 = 3𝑏 1
(V.I.) Answer is (A) 𝟒𝒂 = 𝟑𝒃
The student may read the point (2,9) from the line on the graph.
The student may find the equation 3𝑥 + 𝑦 = 15 joining (5,0) and (0,15) and
17 then verify the point (2,9) satisfies it. 1
Answer is (A) (2,9)
17 Answer is (C) Open Half plane that contains origin, but not the points of the
line 3𝑥 + 5𝑦 = 10 1
(V.I.)
𝟏
Answer is (B) 𝟏𝟎𝟎
The person knows the first 4 digits. So the person has to guess the remaining
18 two digits. 1
1 1 1
P (guessing the PIN )=1×1×1×1×10 × 10 = 100
√3 𝜋 𝜋 𝜋 𝜋 𝜋
sin−1 ( 2 ) + tan−1 1 − sec −1(√2) = 3 + 4 − 4 = 3 ≠ 4
So, A is false.
𝜋 𝜋 𝜋
19 Principal Value branch of sin−1 𝑥 is [− 2 , 2 ] and that of s𝑒𝑐 −1 𝑥 is [0, 𝜋] − {2 }. 1
So, R is true
Answer is (D)Assertion is false, but Reason is true
C. 𝑟⃗ × (𝑎⃗ + 𝑏⃗⃗) = ⃗0⃗ ⇒ 𝑟⃗ is parallel to (𝑎⃗ + 𝑏⃗⃗) and (𝑎⃗ + 𝑏⃗⃗) lies on the plane
of 𝑎⃗ and 𝑏⃗⃗ .
So, 𝑟⃗ is parallel to the plane of 𝑎⃗ and 𝑏⃗⃗ ⇒ 𝑟⃗ is perpendicular to (𝑎⃗ × 𝑏⃗⃗).
20 So, Assertion is true 1
But (𝑎⃗ + 𝑏⃗⃗) lies on the plane of 𝑎⃗ and 𝑏⃗⃗, so (𝑎⃗ + 𝑏⃗⃗) is not perpendicular to
the plane of 𝑎⃗ 𝑎𝑛𝑑 𝑏⃗⃗
Therefore, Reason is false.
Answer is (C) Assertion is true, but Reason is false
Page 3 of 12
Page 5
SECTION B
(VSA type questions of 2 marks each)
21A 𝜋
tan (tan−1(−1) + 3 ) = tan (− 4 + 3 )
𝜋 𝜋 ½
𝜋 𝜋
tan −tan
= 3
𝜋
4
𝜋 1
1+tan tan
3 4
√3−1
= 1+√3 𝑜𝑟 2 − √3 ½
OR OR
21B For domain, −1 ≤ 3𝑥 − 2 ≤ 1 ½
⇒ 1 ≤ 3𝑥 ≤ 3 ½
1 ½
⇒3≤𝑥≤1
1
So, domain of cos−1 (3𝑥 − 2) is [3 , 1] ½
22 𝜋
𝑦 = log tan ( 4 + 2)
𝑥
Differentiating with respect to 𝑥
𝑑𝑦 1 2 𝜋 𝑥 1
= 𝜋 𝑥 .𝑠𝑒𝑐 ( + ) .
½
𝑑𝑥 tan( + ) 4 2 2
4 2
𝜋 𝑥
cos( + ) 1 1
4 2
= 𝜋 𝑥 . 𝜋 𝑥 .2
sin( + ) 𝑐𝑜𝑠2 ( + )
4 2 4 2
1 1 1
= 𝜋 𝑥 𝜋 𝑥 = 𝜋 = cos 𝑥
2 sin( + )cos( + ))
4 2 4 2
sin( +𝑥)
2 1
𝑑𝑦
⟹ 𝑑𝑥 − sec 𝑥 = 0
½
23A (𝑥−3)𝑒 𝑥 (𝑥−1−2)𝑒 𝑥
∫ (𝑥−1)3 𝑑𝑥 = ∫ (𝑥−1)3
𝑑𝑥
1 2 1 𝑑 1
= ∫ ( (𝑥−1)2 − (𝑥−1)3 ) 𝑒 𝑥 𝑑𝑥 = ∫ ((𝑥−1)2 + 𝑑𝑥 ((𝑥−1)2 )) 𝑒 𝑥 𝑑𝑥 1
𝑒𝑥
= (𝑥−1)2 + c (as ∫(𝑓(𝑥) + 𝑓 ′ (𝑥))𝑒 𝑥 𝑑𝑥 = 𝑒 𝑥 𝑓(𝑥) + c) 1
OR OR
4 4
23B A = ∫0 𝑥 𝑑𝑦 = ∫0 √𝑦 𝑑𝑦
2 3 16 1
= 3 × 𝑦 ⁄2 ] 𝑦=4
𝑦=0
= 3 sq. units
1
23B For Visually Impaired:
3 3
A = ∫0 𝑦 𝑑𝑥 = ∫0 √𝑥 𝑑𝑥 1
2 3 1
= 3 × 𝑥 ⁄2 ] 𝑥=3
𝑥=0
= 2√3 sq. units
Page 4 of 12
Page 6
24 Given 𝑓(𝑥 + 𝑦) = 𝑓(𝑥)𝑓(𝑦)
𝑓(0 + 5) = 𝑓(0)𝑓(5)
⟹ 𝑓(0) = 1 ½
′ (5) 𝑓 (5+ℎ)−𝑓(5) 𝑓 (5)𝑓(ℎ)−𝑓(5)
𝑓 = lim = lim [ ∵ 𝑓(𝑥 + 𝑦) = 𝑓(𝑥)𝑓(𝑦 ) ]
ℎ→0 ℎ ℎ→0 ℎ
2𝑓(ℎ)−2
= lim [ ∵ 𝑓(5) = 2] 1
ℎ→0 ℎ
𝑓(ℎ)−1 𝑓(ℎ)−𝑓(0)
= 2 lim = 2 lim = 2 𝑓 ′ (0)
ℎ→0 ℎ ℎ→0 ℎ
= 2 (3) [∵ 𝑓 ′ (0) = 3] ½
=6
25 1
𝑂𝑃 = 2 (4𝑖̂ + 4𝑘̂) = 2𝑖̂ + 2𝑘̂
The vector ⃗⃗⃗⃗⃗⃗ ½
Area of the parallelogram formed by the two adjacent sides as OA and OP
𝑖̂ 𝑗̂ 𝑘̂ ½
⃗⃗⃗⃗⃗⃗ ⃗⃗⃗⃗⃗⃗
= |(𝑂𝐴 X 𝑂𝑃) | = ||1 1 1||
2 0 2
= |2𝑖̂ − 2𝑘̂ | ½
= 2√2 square units. ½
SECTION C
(SA type questions of 3 marks each)
26A 𝑥 𝑦 = 𝑒 𝑥−𝑦
Taking log of both sides
𝑦 log 𝑥 = (𝑥 − 𝑦) log 𝑒
𝑦 log 𝑥 + y = 𝑥 ( since log 𝑒 = 1)
𝑥
⟹ 𝑦 = 1+log 𝑥 1
Differentiating with respect to 𝑥
1
𝑑𝑦 (1+log 𝑥) .1− 𝑥 .
𝑥
= (1+log 𝑥)2
𝑑𝑥
log 𝑥
= (log 𝑒+log 𝑥)2
log 𝑥 1
= (log(𝑥 𝑒))2
𝑑𝑦 log 𝑒 1 1 1
Now ] = = (2𝑙𝑜𝑔𝑒)2 = 22 = 4 (as log 𝑒 = 1) 1
𝑑𝑥 𝑥=𝑒 (log 𝑒 2 )2
Alternative Solution:
𝑥 𝑦 = 𝑒 𝑥−𝑦
Taking log of both sides
𝑦 log 𝑥 = (𝑥 − 𝑦) log 𝑒
𝑦 log 𝑥 + y = 𝑥 ( since log 𝑒 = 1)
Differentiating both sides w.r.t. x
𝑑𝑦 𝑦 𝑑𝑦
log 𝑥 𝑑𝑥 + 𝑥 + 𝑑𝑥 = 1
𝑑𝑦 𝑦
⇒ 𝑑𝑥 (1 + log 𝑥) = 1 − 𝑥
𝑥
𝑑𝑦 𝑥−𝑦 𝑥− 𝑥(1+log 𝑥)−𝑥 𝑥(1+log 𝑥−1) log 𝑥
1+log 𝑥
⇒ 𝑑𝑥 = 𝑥(1+log 𝑥) = 𝑥(1+log 𝑥) = 𝑥(1+log 𝑥)2 = 𝑥(log 𝑒+log 𝑥)2 = (log(𝑥𝑒))2
𝑑𝑦 log 𝑒 1 1 1
Now ] = (log 𝑒 2 )2 = (2𝑙𝑜𝑔𝑒)2 = 22 = 4 (as log 𝑒 = 1)
𝑑𝑥 𝑥=𝑒
Page 5 of 12
Page 7
OR OR
𝑑𝑥 𝑑𝑦
26B = 𝑎(1 − cos 𝜃 ) , 𝑑𝜃 = 𝑎(0 + sin 𝜃), 1
𝑑𝜃
𝑑𝑦
𝑑𝑦 𝑎 sin 𝜃
⇒ 𝑑𝑥 = 𝑑𝜃
𝑑𝑥 = 𝑎(1−cos 𝜃 )
𝑑𝜃
𝜃
2 sin( ) cos( )
𝜃 1
𝜃
= 2
𝜃
2
= cot 2
2𝑠𝑖𝑛2 ( )
2
𝑑2 𝑦 1 𝜃 𝑑𝜃
⇒ 𝑑𝑥 2 = − 2 𝑐𝑜𝑠𝑒𝑐 2 ( 2) 𝑑𝑥
1 𝜃 1
= − 2𝑎 𝑐𝑜𝑠𝑒𝑐 2 ( 2) 𝜃
2𝑠𝑖𝑛2 ( )
2
1 𝜃 1
= − 4𝑎 𝑐𝑜𝑠𝑒𝑐 4 ( 2)
27 Let r be the radius of ice ball at time t .
4
𝑉 = 3 𝜋 𝑟 3 ………… (1)
½
S = 4𝜋𝑟 2 …………..(2)
𝑑𝑉
Given 𝑑𝑡 ∝ S
𝑑𝑉
⟹ 𝑑𝑡 = − k S (where k is some positive constant) ………(3) ½
Differentiating (1) w.r.t. t, we get
𝑑𝑉 4 2 𝑑𝑟
= 𝜋. (3 𝑟 ) 𝑑𝑡
𝑑𝑡 3 1
𝑑𝑉 𝑑𝑟
= 4𝜋𝑟 2 𝑑𝑡 ………. (4)
𝑑𝑡
𝑑𝑟
⟹ - k S = 4𝜋𝑟 2 𝑑𝑡 (from (3) and (4))
½
𝑑𝑟
⟹ - k S = S 𝑑𝑡 (using (2))
⟹ =−k
𝑑𝑟 ½
𝑑𝑡
⇒ radius of the ice-ball decreases at a constant rate
28A
1
2 −1 2
∫−4|𝑥 + 1|𝑑𝑥= ∫−4 (−𝑥 − 1) 𝑑𝑥 + ∫−1(𝑥 + 1) 𝑑𝑥 ½
(𝑥+1)2 −1 (𝑥+1)2 2
= − ] + ] ½
2 −4 2 −1
9 9
= − (0 − 2) + (2 − 0) = 9 ½
It represent the area of shaded region bounded by the curve 𝑦 = |𝑥 + 1|, ½
𝑥 − axis and the lines 𝑥 = −4 𝑎𝑛𝑑 𝑥 = 2
Page 6 of 12
Page 8
OR
28B
1
4 4 𝑥2
Required Area = ∫0 𝑥 𝑑𝑥 − ∫0 4 𝑑𝑥 1
4
𝑥2 1 ½
= 2 ] − 12 [𝑥 3 ]40
0
1 1 16 8 ½
= 2 (16 − 0) − 12 (64 − 0) = 8 − 3 = 3 sq. units
For Visually Impaired:
−𝑥 − 1, 𝑥 < −1
𝑦 = |𝑥 + 1| = 𝑓(𝑥) = {
𝑥 + 1, 𝑥 ≥ −1 1
2 −1 2
∫−4|𝑥 + 1|𝑑𝑥= ∫−4 (−𝑥 − 1) 𝑑𝑥 + ∫−1(𝑥 + 1) 𝑑𝑥
(𝑥+1)2 −1 (𝑥+1)2 2
= − ] + ] 1
2 −4 2 −1
9 9
= − (0 − ) + ( − 0) = 9
2 2
It represent the area of shaded region bounded by the curve 𝑦 = |𝑥 + 1|, 1
𝑥 − axis and the lines 𝑥 = −4 𝑎𝑛𝑑 𝑥 = 2
OR
Page 7 of 12
Page 9
𝑥2 𝑦2 𝑥2 𝑦2 5
25𝑥 2 + 16𝑦 2 = 400 ⇒ 16 + 25 = 1 ⇒ 42 + 52 = 1 ⇒ 𝑦 = 4 √42 − 𝑥 2
45
Required Area = 4 ∫0 4 √42 − 𝑥 2 𝑑𝑥 1
4
𝑥√42 −𝑥 2 42 𝑥
= 5[ + 2 sin−1 (4)] 1
2
0
= 5[0 + 8 sin−1(1) − 0]
𝜋
= 40 × 2 = 20𝜋 sq. units 1
29A The line through (2, −1,3) parallel to the z-axis is given by
𝑟⃗ = (2𝑖̂ − 𝑗̂ + 3𝑘̂ ) + λ(𝑘̂) 1
Any point on this line is P (2, −1, 3 + 𝜆 ) ½
Any point on the given line 𝑟⃗ = (2𝑖̂ − 𝑗̂ + 2𝑘̂) + 𝜇(3𝑖̂ + 6𝑗̂ + 2𝑘̂) is
Q (2 + 3𝜇, −1 + 6𝜇, 2 + 2𝜇)
½
For the intersection point
½
Q (2 + 3𝜇, −1 + 6𝜇, 2 + 2𝜇) = P (2, -1, 3+ λ ) ⇒ 2 = 2 + 3 𝜇 ⇒ 𝜇 = 0
½
The point of intersection is (2, −1,2)
The distance from (2, −1,3) to (2, −1,2) is clearly 1 unit.
Alternative Solution:
Any point on the line through (2, −1, 3) parallel to the z-axis is (2, −1, 𝜆)
Any point on the given line is (2 + 3𝜇, −1 + 6𝜇, 2 + 2𝜇) 1
Therefore, 2 = 2 + 3 𝜇 ⇒ 𝜇 = 0 1
The point of intersection is (2, −1,2) ½
The distance from (2, −1,3) to (2, −1,2) is clearly 1 unit. ½
OR
29B
The line through (2,−1,1)parallel to the z-axis is 𝑟⃗ = (2𝑖̂ − 𝑗̂ + 𝑘̂ ) + λ(𝑘̂) 1
Any point on this line is P (2, −1, 1+ λ )
Any point on the given line is A (3+ 𝜇, 𝜇, 1+ 𝜇 )
A (3+ 𝜇, 𝜇, 1+ 𝜇 ) = P (2, −1, 1+ λ ) ⇒ 𝜇 = −1 1
The point of intersection is (2, −1,0) ½
The distance of (2, −1, 0) from the z-axis is √22 + (−1)2 = √5 units. ½
30 Sketching the graph 1
1
2
Page 8 of 12
Page 10
Corner points A(600,0), B(1200,0), C(800,400), D(400,200)
Values of Z: 𝑍𝐴 = 1200, 𝑍𝐵 = 2400, 𝑍𝐶 = 2000, 𝑍𝐷 = 1000 ½
Maximum 𝑍 = 2400 when 𝑥 = 1200 and 𝑦 = 0 ½
½
For Visually Impaired:
At Corner points A(600,0), B(1200,0), C(800,400), D(400,200)
30 Values of Z are 𝑍𝐴 = 1800, 𝑍𝐵 = 3600, 𝑍𝐶 = 3200, 𝑍𝐷 = 1600 1
Maximum Value of Z = 3600 at B(1200,0) 1
And Minimum Value of Z= 1600 at D(400,200) 1
31 Let the events be:
A: Mehul is selected
B: Rashi is selected
Then according to the question,
A and B are independent events and
𝑃(𝐴) = 0.4, 𝑃(𝐴 ∩ 𝐵̅ ) + 𝑃(𝐵 ∩ 𝐴̅) = 0.5 1
Let 𝑃(𝐵) = 𝑥
Then 𝑃(𝐴 ∩ 𝐵̅ ) + 𝑃(𝐵 ∩ 𝐴̅) = 0.5
⇒ 𝑃(𝐴)𝑃(𝐵̅ ) + 𝑃(𝐵)𝑃(𝐴̅) = 0.5
⇒ 0.4(1 − 𝑥) + 𝑥(1 − 0.4) = 0.5
⇒ 0.4 − 0.4𝑥 + 0.6𝑥 = 0.5
⇒ 0.2𝑥 = 0.5 − 0.4 = 0.1
0.1 1
⇒ 𝑥 = 0.2 = 2 = 0.5 1
So, probability of selection of Rashi = 0.5
Probability of selection of at least one of them = 1 − 𝑃(𝐴̅ ∩ 𝐵̅ )
= 1 − 𝑃(𝐴̅)𝑃(𝐵̅ )
= 1 − 0.6 × 0.5
= 1 − 0.3 = 0.7 1
SECTION D
(LA type questions of 5 marks each)
32 3 −6 −1 1 −2 −1 1 0 0
𝐴𝐵 = [ 2 −5 −1] [0 −1 −1] = [0 1 0] = 𝐼 1
−2 4 1 2 0 3 0 0 1
−1 −1
So, 𝐴 = 𝐵 𝑎𝑛𝑑 𝐵 = 𝐴
½
Given system of equations is
3𝑥 − 6𝑦 − 𝑧 = 3, 2𝑥 − 5𝑦 − 𝑧 + 2 = 0, − 2𝑥 + 4𝑦 + 𝑧 = 5
In matrix form it can be written as: 𝐴𝑋 = 𝐶,
𝑥 ½
3
where 𝑋 = [ ] 𝑎𝑛𝑑 𝐶 = [−2]
𝑦
𝑧 5 ½
Here |𝐴| = −3 − 0 + 2 = −1 ≠ 0 ½
So, the system is consistent and has unique solution given by the
expression 𝑋 = 𝐴−1 𝐶 = 𝐵𝐶 ½
1 −2 −1 3 𝑥 2
⇒ 𝑋 = [0 −1 −1] [−2] ⇒ [𝑦] = [−3] 1
2 0 3 5 𝑧 21
½
Thus 𝑥 = 2, 𝑦 = −3, 𝑧 = 21
Page 9 of 12
Page 11
33A Let 𝑥 = tan 𝜃 ⇒ 𝑑𝑥 = 𝑠𝑒𝑐 2 𝜃 𝑑𝜃 ½
𝜋
log(1+𝑡𝑎𝑛𝜃)
𝐼= ∫ 4
0 1+𝑡𝑎𝑛2 𝜃
. 𝑠𝑒𝑐 2 𝜃 𝑑𝜃
𝜋 𝜋
𝜋
𝐼 = ∫04 𝑙𝑜𝑔 (1 + 𝑡𝑎𝑛𝜃)𝑑𝜃 = ∫0 𝑙𝑜𝑔 [1 + tan (4 − 𝜃)] 𝑑𝜃
4
1
𝜋
1−tan 𝜃
=∫0 log [1 + 1+tan 𝜃] 𝑑𝜃
4
1
𝜋
1+tan 𝜃+1−tan 𝜃
=∫0 log [
4 ] d𝜃
1+tan 𝜃
𝜋
2
=∫04 log [1+tan 𝜃 ] 𝑑𝜃
𝜋 𝜋
=∫0 log 2 𝑑𝜃 − ∫0 log[1 + 𝑡𝑎𝑛𝜃] 𝑑𝜃
4 4
𝜋 1
4
= log 2 × 𝑥]0 − 𝐼
𝜋 1
⇒ 2𝐼 = 4 log 2
𝜋
⇒ 𝐼= log 2
8 ½
33B OR OR
(3 sin 𝜃−2) cos 𝜃 (3 𝑠𝑖𝑛 𝜃−2) 𝑐𝑜𝑠 𝜃
I = ∫ 5−𝑐𝑜𝑠2 𝜃−4 sin 𝜃 𝑑𝜃 = ∫ 5−(1−𝑠𝑖𝑛2 𝜃)−4 𝑠𝑖𝑛 𝜃 𝑑𝜃 ½
𝐿𝑒𝑡 𝑠𝑖𝑛 𝜃 = 𝑡 ⟹ 𝑐𝑜𝑠 𝜃 𝑑𝜃 = 𝑑𝑡
(3𝑡−2)
𝐼 =∫ 5−(1−𝑡 2 )−4𝑡 𝑑𝑡
(3𝑡−2)
1
3𝑡−2
=∫ 𝑡 2 −4𝑡+4 𝑑𝑡 =∫ (𝑡−2)2 𝑑𝑡
3𝑡−2 𝐴 𝐵
Let (𝑡−2)2 = (𝑡−2) + (𝑡−2)2
3𝑡 − 2 = 𝐴(𝑡 − 2) + 𝐵
Comparing the coefficients of t and constant terms on both sides
½
𝐴 = 3 , − 2𝐴 + 𝐵 = − 2 , 𝐵 = 4
(3 sin 𝜃−2) cos 𝜃 3 4
+½
∫ 5−𝑐𝑜𝑠2 𝜃−4 sin 𝜃 𝑑𝜃 =∫ 𝑡−2 𝑑𝑡 + ∫ (𝑡−2)2 dt
4 1+1
= 3log|𝑡 − 2| − 𝑡−2 + 𝐶
4 ½
= 3 log|𝑠𝑖𝑛 𝜃 − 2| − sin 𝜃−2 + C
34A 𝑑
𝑦 + 𝑑𝑥 ( 𝑥𝑦 ) = 𝑥 (sin 𝑥 + 𝑥)
𝑑𝑦
⟹ 𝑦 + (𝑥 𝑑𝑥 + 𝑦 ) = 𝑥 (sin 𝑥 + 𝑥)
𝑑𝑦
1
⟹ 2𝑦 + 𝑥 𝑑𝑥 = 𝑥 (sin 𝑥 + 𝑥)
𝑑𝑦 2𝑦
⟹ 𝑑𝑥 + 𝑥 = (sin 𝑥 + 𝑥)
𝑑𝑦
This a linear differential equation of the form 𝑑𝑥 + 𝑃𝑦 = 𝑄
2
P= 𝑥 ,Q=(sin 𝑥 + 𝑥)
2
1
∫ 𝑥 dx 2 𝑙𝑜𝑔 𝑥 𝑙𝑜𝑔 𝑥 2 2
I.F =𝑒 = 𝑒 =𝑒 =𝑥
Solution will be 𝑦 . I.F =∫ 𝑄. 𝐼𝐹 𝑑𝑥 1
𝑦𝑥 2 = ∫(𝑠𝑖𝑛 𝑥 + 𝑥) 𝑥 2 dx
𝑦𝑥 2 =∫ 𝑠𝑖𝑛𝑥. 𝑥 2 𝑑𝑥 + ∫ 𝑥 3 𝑑𝑥
Page 10 of 12
Page 12
𝑥4 1
⟹ 𝑦𝑥 2 = −𝑥 2 cos 𝑥 + 2 ∫ 𝑥 cos 𝑥 𝑑𝑥 + 4 + C
𝑥4
⟹ 𝑦𝑥 2 = −𝑥 2 𝑐𝑜𝑠 𝑥 +2 (𝑥 𝑠𝑖𝑛𝑥 + 𝑐𝑜𝑠 𝑥) + 4 + C
Which is the required solution 1
OR
𝑥 𝑥
2𝑦 𝑒 ⁄𝑦 𝑑𝑥 + ( 𝑦 − 2𝑥 𝑒 ⁄𝑦 ) 𝑑𝑦 = 0
34B 𝑥 𝑥 𝑦
𝑥
𝑑𝑥 2𝑥 𝑒 ⁄𝑦 −𝑦 2 𝑒 −1
𝑦
⟹ = 𝑥 = 𝑥
𝑑𝑦 2𝑦 𝑒 ⁄𝑦 2 𝑒𝑦 1
It is a homogeneous differential equation.
𝑑𝑥 𝑑𝑣
Let 𝑥 = 𝑣𝑦 ⟹ 𝑑𝑦 = 𝑣 +𝑦 𝑑𝑦 1
𝑑𝑣 2𝑣𝑒 𝑣 −1
𝑣 +𝑦 𝑑𝑦 = 2𝑒 𝑣
𝑑𝑣 2𝑣𝑒 𝑣 −1 2𝑣𝑒 𝑣 −1−2𝑣𝑒 𝑣
⟹ 𝑦 𝑑𝑦 = –𝑣=
2𝑒 𝑣 2𝑒 𝑣
𝑑𝑣 −1
⟹ 𝑦 =
𝑑𝑦 2𝑒 𝑣
𝑣 𝑑𝑦
⟹ 2𝑒 𝑑𝑣 = − 𝑦
1
𝑣 𝑑𝑦
∫ 2𝑒 𝑑𝑣 = − ∫ 𝑦
⟹ 2𝑒 𝑣 = − log |𝑦| + C
𝑥
⟹ 2𝑒 𝑦 +" 𝑙𝑜𝑔|𝑦| = C 1
When 𝑥 = 0 , 𝑦 = 1 , C = 2
𝑥
Required solution 2 𝑒 𝑦 + log|𝑦| = 2
1
35 Let
𝑥−1 𝑦−0
= −1 =
𝑧+1
=𝜆 ⇒ Any point on it is (3 𝜆 +1, − 𝜆, −1) ½
3 0
1
For the point where 𝑦 = 1 ⇒ 𝜆 = −1
½
⇒ The point is (−2, 1 , −1)
1
The directions of the two lines are 𝑚 ⃗⃗⃗ = 3𝑖̂ − 𝑗̂
½
and 𝑛⃗⃗ = −2𝑖̂ + 2𝑗̂ + 𝑘̂
𝑖̂ 𝑗̂ 𝑘̂
⃗⃗⃗ × 𝑛⃗⃗ = | 3 −1 0| = −𝑖̂ − 3𝑗̂ + 4𝑘̂
𝑚 1
−2 2 1
The equation of the required line is
𝑟⃗ = (−2𝑖̂ + 𝑗̂ − 𝑘̂) + 𝜇(−𝑖̂ − 3𝑗̂ + 4𝑘̂) ½
Alternative Solution:
𝑥−1 𝑦−0 𝑧+1
Let = −1 = = 𝜆 ⇒ Any point on it is (3 𝜆 + 1, − 𝜆, −1) ½
3 0
For the point where 𝑦 = 1 ⇒ 𝜆 = −1 1
⇒ The point is (−2, 1, −1) ½
Let the direction ratios of the required line be a, b, c
Then 3𝑎 − 𝑏 = 0
And −2𝑎 + 2𝑏 + 𝑐 = 0 1
𝑎 −𝑏 𝑐 𝑎 𝑏 𝑐 1
Solving we get = 3 = 4 ⇒ −1 = −3 = 4
−1
𝑥+2 𝑦−1 𝑧+1
The required line is = −3 = =𝜇 ½
−1 4
In vector form 𝑟⃗ = (−2𝑖̂ + 𝑗̂ − 𝑘̂) + 𝜇(−𝑖̂ − 3𝑗̂ + 4𝑘̂) ½
Page 11 of 12
Page 13
SECTION- E
(3 case-study/passage-based questions of 4 marks each)
36 I. Traffic flow is not reflexive as (𝐴, 𝐴) ∉ 𝑅 (or no major spot is connected
with itself) 1
II. Traffic flow is not transitive as (𝐴, 𝐵) ∈ 𝑅 𝑎𝑛𝑑 (𝐵, 𝐸) ∈ 𝑅, 𝑏𝑢𝑡 (𝐴, 𝐸) ∉ 𝑅 1
III A. 𝑅 = {(𝐴, 𝐵), (𝐴, 𝐶), (𝐴, 𝐷), (𝐵, 𝐶), (𝐵, 𝐸), (𝐶, 𝐸), (𝐷, 𝐸), (𝐷, 𝐶)} 1
Domain = {𝐴, 𝐵, 𝐶, 𝐷} ½+
Range = {𝐵, 𝐶, 𝐷, 𝐸} ½
OR
III B. No, the traffic flow doesn’t represent a function as A has three images. 1+1
37 I. 𝑃(𝑥) = 𝑅(𝑥) − 𝐶(𝑥) = −0.3𝑥 2 + 20𝑥 − (0.5𝑥 2 − 10𝑥 + 150)
= −0.8𝑥 2 + 30𝑥 − 150 1
II. For critical points 𝑃′ (𝑥) = 0 ⇒ −1.6𝑥 + 30 = 0
30 300 1
⇒ 𝑥 = 1.6 = 16 = 18.75
III A. Now 𝑃′′ (𝑥) = −1.6
1
In particular 𝑃′′ (18.75) = −1.6 < 0
1
So, critical value 𝑥 = 18.75 corresponds to a maximum profit.
OR
III B. As 𝑥 is the number of bulbs, so practically 18 bulbs correspond to a
maximum profit. 1
Maximum profit is 𝑃(18) = −0.8 × 182 + 30 × 18 − 150
= −259.2 + 540 − 150 1
= 540 − 409.2 = ₹130.80
38 Let the events be
E1: the student is in the first group (time spent on screen is more than 4 hours)
E2: the student is in the second group (time spent on screen is 2 to 4 hours)
E3: the student is in third group (time spent on screen is less than 2 hours)
A: the event of the student showing symptoms of anxiety and low retention
60 30 10
P(E1) = 100 P(E2) = 100 and P(E3) = 100
80 70 30
P(A|E1) = 100 P(A|E2) = 100 and P(A|E3) = 100
I. P(A)= P(E1) x P(A|E1) + P(E2) x P(A|E2) + P(E3) x P(A|E3)
60 80 30 70 10 30 72
= 100 x 100 + 100 x 100 + 100 x 100 = 100 = 72% 2
𝑃(𝐸1 ∩𝐴)
II. P(E1|A) = 𝑃(𝐴)
60 80
( x
100 100
) 48 2 2
= 72 = 72 = 3
( )
100
Page 12 of 12
Page 14
गणित – 041
अंक योजना
कक्षा - बारहव ं (2025 - 26)
क्रम उत्तर संकेत /मूल्य बिन्दु अंक
संख्या बिभाजन
−1 1
ग्राफ से स्पष्ट है कक प्ांत [ 2 , 2] है
1 तो आलेख फलन sin−1 (2𝑥) का है 1
उत्तर है (B) 𝐬𝐢𝐧−1 (2𝒙)
प्ांत
[ 3 , 3] है ,
−1 1
1
(V.I.) अत : फलन cos −1 (3𝑥) है । 1
उत्तर है (C) 𝐜𝐨𝐬 −1 (3𝒙)
AB पररभाबित है इसबलए n=4
AC पररभाबित है इसबलए p=4
2 AB तथा AC एक ही कोरि के िर्ग आव्यूह हैं । 1
इसबलए 𝑚 × 3 = 𝑚 × 𝑞 ⇒ 𝑞 = 3 = 𝑚
उत्तर (A) है 𝒎 = 𝒒 = 3तथा𝒏 = 𝒑 = 4
क्योंकक A एक बििम समबमत है
इसबलए 𝑝 = 0, 𝑞 = 2, 𝑟 = −3, 𝑡 = 4
𝑞+𝑡 6
3 इसबलए = = −2 1
𝑝+𝑟 −3
उत्तर है (A) -2
𝑎𝑑𝑗𝐴 = 27⇒ |𝐴|3 = 27 = 33 ⇒ |𝐴| = 3
𝐴(𝑎𝑑𝑗𝐴) = |𝐴|𝐼 = 3𝐼
4 1
उत्तर है (C) 3𝑰
1
0 0
3 0 0 3
आव्यूह[0 0] का व्युत्क्क्रम=[
1
2 0 0]
2
5 0 0 5 1 1
0 0
5
अत : उत्तर (B) है.
cos67𝑜 sin67𝑜
| | = cos67𝑜 cos23𝑜 − sin67𝑜 sin23𝑜 = cos(67𝑜 + 23𝑜 ) = cos90𝑜 = 0
sin23𝑜 cos23𝑜
6 1
उत्तर है (A) 0
𝑓(𝑥) , 𝑥 = 𝜋 पर संतत है
⇒ lim−1 (𝑘𝑥 + 1)= lim+1 𝑐𝑜𝑠𝑥=f(π)
𝑥→𝜋 𝑥→𝜋
⇒ lim [𝑘(𝜋 − ℎ) + 1] = lim cos(𝜋 + ℎ) = 𝑘𝜋 + 1
ℎ→0 ℎ→0
7 ⇒ 𝑘𝜋 + 1 = −1 1
−2
⇒𝑘= 𝜋
−2
उत्तर (D) 𝝅
पेज 1 of 12
Page 15
𝑓(𝑥) = 𝑥tan−1 𝑥
1
𝑓 ′ (𝑥) = 1. tan−1 𝑥 + 𝑥. 1+𝑥2
8 𝑓 ′ (1) =1 . 1tan−1 +
1 𝜋 1
= + 1
1+1 4 2
𝝅 1
उत्तर है (B) 4 + 2
𝑓(𝑥) = 10 − 𝑥 − 2𝑥 2
⇒ 𝑓 ′ (𝑥) = −1 − 4𝑥
िर्गमान फलन के बलए 𝑓 ′ (𝑥) ≥ 0
⇒ −(1 + 4𝑥) ≥ 0
⇒ (1 + 4𝑥) ≤ 0
1
9 ⇒𝑥≤− 1
4
1
⇒ 𝑥 ∈ (−∞, − ]
4
1
उत्तर है (A) (−∞, − 4]
1
उत्तर (A) (−∞, − 4]
𝑥𝑑𝑥 + 𝑦𝑑𝑦 = 0
⇒ ∫ 𝑥𝑑𝑥 = −∫ 𝑦𝑑𝑦
𝑥2 −𝑦 2
⇒ = +𝑘
2 2
10 2 2 1
⇒ 𝑥 + 𝑦 = 2𝑘
अतः हल है ⇒ 𝑥 2 + 𝑦 2 = 2𝑘 , k एक स्िेच्छ अचर है
उत्तर है (C) ,िृत
𝑏 𝑏 𝑏
𝐼 =∫𝑎 𝑥𝑓(𝑥)𝑑𝑥 = ∫𝑎 (𝑎 + 𝑏 − 𝑥)𝑓(𝑎 + 𝑏 − 𝑥)𝑑𝑥 ⇒ 𝐼 = ∫𝑎 (𝑎 + 𝑏 − 𝑥)𝑓(𝑥)𝑑𝑥 (given
𝑓(𝑎 + 𝑏 − 𝑥) = 𝑓(𝑥) )
𝑏 𝑏
⇒ 𝐼 ∫𝑎 (𝑎 + 𝑏) 𝑓(𝑥)𝑑𝑥 - ∫𝑎 𝑥𝑓(𝑥)𝑑𝑥
𝑏
11 ⇒2 𝐼= (𝑎 + 𝑏) ∫𝑎 𝑓(𝑥)𝑑𝑥 1
1 𝑏
⇒ 𝐼 = 2 (𝑎 + 𝑏) ∫𝑎 𝑓(𝑥)𝑑𝑥
𝒂+𝒃 𝒃
उत्तर है (D) 2 ∫𝒂 𝒇(𝒙)𝒅𝒙
माना 𝐼 =∫ 𝑥 3 sin4 (𝑥 4 )cos(𝑥 4 ) 𝑑𝑥माना sin(𝑥 4 ) = 𝑡 ⇒ 4𝑥 3 cos(𝑥 4 ) 𝑑𝑥 = 𝑑𝑡 ⇒
1
𝑥 3 cos(𝑥 4 ) = 4 𝑑𝑡
1 1 5 1
ति 𝐼 = ∫ 𝑡 4 ( 𝑑𝑡) = 𝑡 + C = sin5(𝑥 4 ) + C
4 20 20
12 1
⇒ 𝐼 = 20 sin5 (𝑥 4 ) + 𝐶 = 𝑎 sin (𝑥 ) + C 5 4 1
1
अत : 𝑎 = 20
1
उत्तर (B)
20
1𝑥1+2𝑥2+1𝑥3
सकदश आव्यूह 𝑖̂ + 2𝑗̂ + 𝑘̂ का रे खा 𝑟⃗ = (3𝑖̂ − 𝑗̂) + 𝜆(𝑖̂ + 2𝑗̂ + 3𝑘̂ ) पर प्क्षेप है =
√12 +22 +32
8
इकाई
13 √14 1
8
उत्तर है (C) √14 इकाई
पेज 2 of 12
Page 16
बिंद ु (a, b, c) की y-अक्ष से दूरी √𝑎2 + 𝑐 2 है
14 तो, दूरी =√32 + 52 =√34 इकाई है। 1
उत्तर है (B) √34 इकाई
(2𝑎⃗. 𝑖̂)𝑖̂ − (𝑏⃗⃗. 𝑗̂)𝑗̂ + (𝑐⃗. 𝑘̂ )𝑘̂ = (2 × 3)𝑖̂ − (1)𝑗̂ + (2)𝑘̂
= 6𝑖̂ − 𝑗̂ + 2𝑘̂ = 𝑐⃗
15 1
उत्तर है (D) 𝒄⃗⃗
बिंद ु (1,0) और (0,2) समीकरण 2𝑥 + 𝑦 = 2को संतुष्ट करते हैं
और छायांककत क्षेत्र से पता चलता है कक (0,0) सुसंर्त क्षेत्र में बस्थत नहीं है
16 तो, असबमका 2𝑥 + 𝑦 ≥ 2 है 1
उत्तर है (B) 2𝒙 + 𝒚 ≥ 2
(4,0) और (0,3) अबर्कतम मान देते हैं इसबलए
16 𝑍(4,0) = 𝑍(0,3) ⇒ 4𝑎 + 𝑐 = 3𝑏 + 𝑐 ⇒ 4𝑎 = 3𝑏
1
(V.I.) उत्तर है (A) 4𝒂 = 3𝒃
छात्र ग्राफ पर िनी रे खा से बिंद ु (2,9) को पढ़ सकता है।
बिन्दु (5,0) और (0,15) को बमलाने पर छात्र समीकरण 3𝑥 + 𝑦 = 15 प्ाप्त कर सकता है
17 और कफर बिंद ु (2,9) को सत्क्याबपत कर सकता है जो इसे संतुष्ट करता है। 1
उत्तर है (A) (2,9)
उत्तर है (C) खुला अर्ग तल बजसमें मूल बिंद ु शाबमल है, लेककन रे खा 3𝑥 + 5𝑦 = 10 के
17-VI बिंद ु नहीं हैं 1
1
उत्तर है (B) 100
व्यबि को पहले 4 अंक पता हैं, इसबलए व्यबि को शेि दो अंकों का अनुमान लर्ाना होर्ा।
18 1
P (बपन का अनुमान लर्ाना) =1×1×1×1×10 × 10 = 100
1 1 1
√3 𝜋 𝜋 𝜋 𝜋 𝜋
sin−1 ( 2 ) + tan−1 1 − 𝑠𝑒𝑐 −1 (√2) = 3 + 4 − 4 = 3 ≠ 4
अत :A असत्क्य है
sin−1 𝑥 की मुख्य शाखा 𝑖𝑠[ 2 , 2 ] है तथा 𝑠𝑒𝑐 −1 𝑥की[0, 𝜋] − { 2 }
−𝜋 𝜋 𝜋
19 1
अतः , R सत्क्य है
उत्तर (D) है , अबभकथन असत्क्य तथा तकग सत्क्य
𝑟⃗ × (𝑎⃗ + 𝑏⃗⃗) = 0
⃗⃗
=>𝑟⃗ ,(𝑎⃗ + 𝑏⃗⃗) के समांतर है तथा (𝑎⃗ + 𝑏⃗⃗) , 𝑎⃗ और 𝑏⃗⃗ के तल पर बस्थत है, इसबलए 𝑟⃗,
𝑎⃗, 𝑏⃗⃗ तल के समांतर है ⇒ 𝑟⃗ , (𝑎⃗ × 𝑏⃗⃗).के लंिित है अतः अबभकथन सत्क्य है
20 1
लेककन(𝑎⃗ + 𝑏⃗⃗), 𝑎⃗ और 𝑏⃗⃗ के तल पर बस्थत है, इसबलए (𝑎⃗ + 𝑏⃗⃗) , 𝑎⃗ और𝑏⃗⃗ के तल पर
लंिित नहीं है इसबलए, तकग र्लत है।
उत्तर है (C) अबभकथन सत्क्य है, लेककन तकग र्लत है
पेज 3 of 12
Page 17
खंड - ख ,(VSA)प्कार के प्श्न हैं ,
प्त्क्यक
े के 2 अंक हैं
𝜋 −𝜋 𝜋
21A tan(tan−1 (−1) + ) = tan( + ) ½
3 4 3
𝜋 𝜋
tan −tan
3 4
= 1+tan𝜋
tan
𝜋 1
3 4
3−1
=√1+ 3 = 2 − √3
√ ½
अथिा
अथिा
प्ांत के बलए , −1 ≤ 3𝑥 − 2 ≤ 1
21B
⇒ 1 ≤ 3𝑥 ≤ 3 ½
1
⇒3≤𝑥≤1 ½
1 ½
अतः cos−1 (3𝑥 − 2)का प्ांत है [ , 1]
3 ½
𝜋 𝑥
22 𝑦 = logtan( 4 + 2)
x के सापेक्ष अिकलन करने पर
𝑑𝑦 1 2 𝜋 𝑥 1
= 𝜋 𝑥 .𝑠𝑒𝑐 ( + ). ½
𝑑𝑥 tan( + ) 4 2 2
4 2
𝜋 𝑥
cos( + ) 1 1
4 2
= 𝜋 𝑥 . 𝜋 𝑥 .
sin( + ) cos2 ( + ) 2
4 2 4 2
1
= 2sin(𝜋+𝑥)cos(𝜋+𝑥)) = sin(𝜋+𝑥) = cos𝑥
1 1 1
4 2 4 2 2
𝑑𝑦
=⟹ 𝑑𝑥 − 𝑠𝑒𝑐𝑥 = 0 ½
23A (𝑥−3)𝑒 𝑥 (𝑥−1−2)𝑒 𝑥 1
∫ 3 𝑑𝑥 ∫ 𝑑𝑥
(𝑥−1) (𝑥−1)3
1 2 1 𝑑 1
∫ ((𝑥−1)2 − (𝑥−1)3 )𝑒 𝑥 𝑑𝑥 = ∫ ((𝑥−1)2 + 𝑑𝑥 ((𝑥−1)2 ))𝑒 𝑥 𝑑𝑥
𝑒𝑥 1
(𝑥−1)2
+c (क्योंकक ∫ (𝑓(𝑥) + 𝑓 ′ (𝑥))𝑒 𝑥 𝑑𝑥 = 𝑒 𝑥 𝑓(𝑥) + c)
अथिा अथिा
4 4
A = ∫0 𝑥𝑑𝑦 = ∫0 √𝑦 𝑑𝑦
23B 1
3
= 3 िर्ग इकाई
2 𝑦=4 16
= [3 × 𝑦 2 ]𝑦=0 1
23B दृबष्टिाबर्तों के बलए:
3
A = ∫0 𝑦𝑑𝑥 =
3
∫0 √𝑥 𝑑𝑥 1
3 1
= 2√3 िर्ग इकाई
2 𝑥=3
=[3 × 𝑥 2 ]𝑥=0
24 कदया है 𝑓(𝑥 + 𝑦) = 𝑓(𝑥)𝑓(𝑦)
𝑓(0 + 5) = 𝑓(0)𝑓(5)
⟹ 𝑓(0) = 1 ½
′ 𝑓(5+ℎ)−𝑓(5) 𝑓(5)𝑓(ℎ)−𝑓(5)
𝑓 (5) = lim ℎ
= lim ℎ
[∵ 𝑓(𝑥 + 𝑦) = 𝑓(𝑥)𝑓(𝑦)]
ℎ→0 ℎ→0
2𝑓(ℎ)−2
= lim [∵ 𝑓(5) = 2]
ℎ→0 ℎ 1
𝑓(ℎ)−1 𝑓(ℎ)−𝑓(0)
=2 lim
ℎ
= 2lim ℎ
=2 𝑓 ′ (0)
ℎ→0 ℎ→0
= 2 (3) [∵ 𝑓 ′ (0) = 3]
=6 ½
पेज 4 of 12
Page 18
25 सकदश ⃗⃗⃗⃗⃗⃗ = 1 (4𝑖̂ + 4𝑘̂ ) = 2𝑖̂ + 2𝑘̂
𝑂𝑃 ½
2
दो आसन्न भुजाओं OA और OP द्वारा बनर्मगत समांतर चतुभुगज का क्षेत्रफल
𝑖̂ 𝑗̂ 𝑘̂
⃗⃗⃗⃗⃗⃗ × 𝑂𝑃
=|(𝑂𝐴 ⃗⃗⃗⃗⃗⃗)|=|1 1 1| ½
2 0 2
=|2𝑖̂ − 2𝑘̂ | ½
=2√2 िर्ग इकाई। ½
खंड - र् लघु उत्तर ( SA ) प्कार के प्श्न,
प्त्क्यक
े के 3 अंक
26A 𝑥 𝑦 = 𝑒 𝑥−𝑦
दोनों पक्षों का log लेने पर
𝑦log𝑥 = (𝑥 − 𝑦)log𝑒
𝑦log𝑥 + y = 𝑥 ( क्योंकक log𝑒 = 1)
𝑥
⟹ 𝑦=
1+log𝑥
1
𝑥 के सापेक्ष अिकलन करने पर
1
𝑑𝑦 (1+log𝑥).1−𝑥.
𝑥
𝑑𝑥
= (1+log𝑥)2
log𝑥
=(log𝑒+log𝑥)2 1
log𝑥
=(log(𝑥𝑒))2
𝑑𝑦 log𝑒 1 1 1 1
Now 𝑑𝑥 ]𝑥=𝑒 = (log𝑒 2 )2 = (2𝑙𝑜𝑔𝑒)2 = 22= 4 (∵ log𝑒 = 1)
िैकबल्पक हल :
𝑥 𝑦 = 𝑒 𝑥−𝑦
दोनों पक्षों का log लेने पर
𝑦log𝑥 = (𝑥 − 𝑦)log𝑒
𝑦log𝑥 + y = 𝑥 ( since log𝑒 = 1)
x के सापेक्ष अिकलन करने पर
𝑑𝑦 𝑦 𝑑𝑦
log𝑥 𝑑𝑥 + 𝑥 + 𝑑𝑥 = 1
𝑑𝑦 𝑦
⇒ 𝑑𝑥 (1 + log𝑥) = 1 − 𝑥
𝑥
𝑑𝑦 𝑥−𝑦 𝑥− 𝑥(1+log𝑥)−𝑥 𝑥(1+log𝑥−1) log𝑥
1+log𝑥
⇒ 𝑑𝑥 = 𝑥(1+log𝑥) = 𝑥(1+log𝑥) = 𝑥(1+log𝑥)2 = 𝑥(log𝑒+log𝑥)2 = (log(𝑥𝑒))2
1 1 1
अि
𝑑𝑦 log𝑒
] =
𝑑𝑥 𝑥=𝑒 (log𝑒 2 )2
= (2𝑙𝑜𝑔𝑒)2 = 2= 4 (∵ log𝑒 = 1)
2
अथिा
𝑑𝑥 𝑑𝑦
= 𝑎(1 − cos𝜃) , 𝑑𝜃 = 𝑎(0 + sin𝜃),
𝑑𝜃
𝑑𝑦
अथिा
26B 𝑑𝑦
⇒ 𝑑𝑥 = 𝑑𝜃
𝑎sin𝜃
𝑑𝑥 =𝑎(1−cos𝜃)
𝑑𝜃
𝜃 𝜃
2sin( )cos( ) 𝜃 1
2 2
𝜃 =cot 2
2sin2 ( )
2
𝑑2 𝑦 1 𝜃 𝑑𝜃 1
⇒
𝑑𝑥 2
=− 2 𝑐𝑜𝑠𝑒𝑐 2 (2 ) 𝑑𝑥
−1 𝜃 1
= 2𝑎 𝑐𝑜𝑠𝑒𝑐 2 (2 ) 2 𝜃
2sin ( )
2
1 4 𝜃
= − 4𝑎 𝑐𝑜𝑠𝑒𝑐 (2 )
1
पेज 5 of 12
Page 19
27 मान लीबजए समय t पर िफग की र्ेंद की बत्रज्या r है
4
𝑉 = 3 𝜋𝑟 3 ………… (1)
½
S = 4𝜋𝑟 2 …………..(2)
कदया है 𝑑𝑡 ∝ 𝑆
𝑑𝑉
=− k S (जहााँ k कोई र्नात्क्मक बस्थरांक है) ------- (3)
𝑑𝑉
⟹ ½
𝑑𝑡
t के सापेक्ष (1) का अिकलन करने पर हमें प्ाप्त होता है
𝑑𝑉 4 𝑑𝑟
= 3 𝜋.(3 𝑟 2 ) 𝑑𝑡
𝑑𝑡
𝑑𝑉 𝑑𝑟
1
𝑑𝑡
=4𝜋𝑟 2 𝑑𝑡 ………. (4)
(3) and (4) द्वारा
𝑑𝑟
⟹ - k S = 4𝜋𝑟 2 𝑑𝑡 ½
( (2) का इस्तेमाल करने पर )
𝑑𝑟
⟹ - k S = S 𝑑𝑡
𝑑𝑟
⟹ = −k ½
𝑑𝑡
⇒िफग के र्ोले की बत्रज्या बनबित दर से घि रही है
28A
1
2 −1 2
∫−4 |𝑥 + 1|𝑑𝑥 = ∫−4 (−𝑥 − 1) 𝑑𝑥 + ∫−1(𝑥 + 1) 𝑑𝑥 ½
−(𝑥+1)2 −1 (𝑥+1)2 2 ½
= 2
]−4 +
2
]−1
9 9 ½
= −(0 − 2) + (2 − 0) = 9
½
यह िक्र 𝑦 = |𝑥 + 1|,x अक्ष तथा रे खाओं 𝑥 =−4 तथा 𝑥 =2 द्वारा छायांककत क्षेत्र से बघरा हुआ क्षेत्रफल
दशागता है
28B अथिा
1
पेज 6 of 12
Page 20
4 4 𝑥2 1
आिश्यक क्षेत्र = ∫0 𝑥𝑑𝑥 − ∫0 4 𝑑𝑥
𝑥2 4 1
= ] − 12 ½
2 0
1 1 16 8 ½
= 2 (16 − 0) − 12 (64 − 0) = 8 − 3 = 3 िर्ग इकाई
दृबष्टिाबर्तों के बलए:
−𝑥 − 1, 𝑥 < −1
𝑦 = |𝑥 + 1| = 𝑓(𝑥) = {
𝑥 + 1, 𝑥 ≥ −1 1
2 −1 2
∫−4 |𝑥 + 1|𝑑𝑥 = ∫−4 (−𝑥 − 1) 𝑑𝑥 + ∫−1(𝑥 + 1) 𝑑𝑥
=
−(𝑥+1)2 −1
] +
(𝑥+1)2 2
]−1 1
2 −4 2
9 9
= −(0 − 2) + (2 − 0) = 9
1
यह िक्र 𝑦 = |𝑥 + 1|,x अक्ष तथा रे खाओं 𝑥 =−4 तथा 𝑥 =2 द्वारा छायांककत क्षेत्र से बघरा हुआ
क्षेत्रफल दशागता है ।
अथिा
𝑥2 𝑦2 𝑥2 𝑦2 5
25𝑥 2 + 16𝑦 2 = 400 ⇒ 16 + 25 = 1 ⇒ 42 + 52 = 1 ⇒ 𝑦 = 4 √42 − 𝑥 2
45
आिश्यक क्षेत्र = 4 ∫0 4 √42 − 𝑥 2 𝑑𝑥
1
𝑥√4 2 −𝑥 2 42 𝑥
= 5[ + sin−1 ( )]40 1
2 2 4
= 5[0 + 8sin−1 (1) − 0]
1
𝜋
= 40 × 2 = 20𝜋 िर्ग इकाई
29A z-अक्ष के समांतर (2, −1,3)से र्ुजरने िाली रे खा 𝑟⃗ = (2𝑖̂ − 𝑗̂ + 3𝑘̂ ) + 𝜆(𝑘̂ ) द्वारा दी र्ई है
इस रे खा पर कोई बिंद ु P (2,−1,3+ 𝜆)है 1
दी र्ई रे खा 𝑟⃗ = (2𝑖̂ − 𝑗̂ + 2𝑘̂ ) + 𝜇(3𝑖̂ + 6𝑗̂ + 2𝑘̂ )पर कोई बिंद ु ½
Q (2 + 3𝜇, −1 + 6𝜇, 2 + 2𝜇)है
पेज 7 of 12
Page 21
प्बतच्छेदन बिंद ु के बलए
Q (2 + 3𝜇, −1 + 6𝜇, 2 + 2𝜇) = P (2, -1, 3+ λ ) ⇒ 2 = 2 + 3 𝜇 ⇒ 𝜇 = 0 है ½
½
प्बतच्छेदन बिंद ु है (2,−1,2)
(2,−1,3)की (2,−1,2) से दूरी स्पष्ट रूप से 1 इकाई है। ½
िैकबल्पक हल :
बिन्दु (2,−1,3) से z अक्ष के समांतर रे खा पर कोई बिन्दु (2,−1, 𝜆) है 1
दी र्ई रे खा पर कोई बिन्दु (2 + 3𝜇, −1 + 6𝜇, 2 + 2𝜇)है
इसबलए 2 = 2 + 3 𝜇 ⇒ 𝜇 = 0 1
प्बतच्छेदन बिंद ु है (2,−1,2)
½
(2,−1,3)की (2,−1,2) से दूरी स्पष्ट रूप से 1 इकाई है।
½
अथिा
29B
(2,−1,1) से होकर जाने िाली z-अक्ष के समांतर रेखा है 𝑟⃗ = (2𝑖̂ − 𝑗̂ + 𝑘̂ ) + λ(𝑘̂ ) इस 1
रे खा पर कोई बिंद ु P (2, −1, 1+ λ) है
दी र्ई रे खा पर कोई बिंद ु A (3+ 𝜇, 𝜇, 1+ 𝜇 ) है
A (3+ 𝜇, 𝜇, 1+ 𝜇 ) = P (2, −1, 1+ λ ) ⇒ 𝜇 = −1 1
प्बतच्छेद बिंद ु (2, −1,0) है ½
z-अक्ष से (2,−1,0) की दूरी है √22 + (−1)2 =√5 इकाई ½
30 ग्राफ का रे खाबचत्र िनाना 1
1
2
30 कोनीय बिंद ु A(600,0), B(1200,0), C(800,400), D(400,200) हैं ।
½
½
Z के मान:𝑍𝐴 = 1200, 𝑍𝐵 = 2400, 𝑍𝐶 = 2000, 𝑍𝐷 = 1000 ½
अबर्कतम 𝑍 = 2400जि 𝑥 = 1200और𝑦 = 0
दृबष्टिाबर्त लोर्ों के बलए:
कोनीय बिंद ु A(600,0), B(1200,0), C(800,400), D(400,200) पर 1
Z के मान हैं 𝑍𝐴 = 1800, 𝑍𝐵 = 3600, 𝑍𝐶 = 3200, 𝑍𝐷 = 1600
1
B(1200,0) पर Z का अबर्कतम मान = 3600
और D(400,200) पर Z का न्यूनतम मान = 1600 1
पेज 8 of 12
Page 22
31 मान लीबजए घिनाएाँ हैं:
A: मेहुल का चयन ककया र्या है
B: राबश का चयन ककया र्या है
तो प्श्न के अनुसार,
A और B स्ितंत्र घिनाएाँ हैं और
1
𝑃(𝐴) = 0.4, 𝑃(𝐴 ∩ 𝐵̅) + 𝑃(𝐵 ∩ 𝐴̅) = 0.5
माना 𝑃(𝐵) = 𝑥
तो 𝑃(𝐴 ∩ 𝐵̅) + 𝑃(𝐵 ∩ 𝐴̅) = 0.5
⇒ 𝑃(𝐴)𝑃(𝐵̅) + 𝑃(𝐵)𝑃(𝐴̅) = 0.5
⇒ 0.4(1 − 𝑥) + 𝑥(1 − 0.4) = 0.5
⇒ 0.4 − 0.4𝑥 + 0.6𝑥 = 0.5
⇒ 0.2𝑥 = 0.5 − 0.4 = 0.1
0.1 1
⇒𝑥= = = 0.5
0.2 2 1
तो, राबश के चयन की संभािना = 0.5
उनमें से कम से कम एक के चयन की संभािना 1 − 𝑃(𝐴̅ ∩ 𝐵̅)
1 − 𝑃(𝐴̅)𝑃(𝐵̅)
1 − 0.6 × 0.5
1
1 − 0.3 = 0.7
खंड घ दीर्ग उत्तर (LA)प्कार के प्श्न हैं ,
प्त्क्यक
े 5 अंक का है )
32 3 −6 −1 1 −2 −1 1 0 0
𝐴𝐵 = [ 2 −5 −1][0 −1 −1] = [0 1 0] = 𝐼 1
−2 4 1 2 0 3 0 0 1
इसबलए,𝐴−1 = 𝐵तथा𝐵−1 = 𝐴
समीकरणों का कदया हुआ बनकाय ½
3𝑥 − 6𝑦 − 𝑧 = 3,2𝑥 − 5𝑦 − 𝑧 + 2 = 0, −2𝑥 + 4𝑦 + 𝑧 = 5
आव्यूह रूप में इसे इस प्कार बलखा जा सकता है: 𝐴𝑋 = 𝐶 , ½
𝑥 3
जहााँ 𝑋 = [𝑦]तथा𝐶 = [−2] ½
𝑧 5
यहााँ |𝐴| = −3 − 0 + 2 = −1 ≠ 0 ½
इसबलए,बनकाय सुसंर्त है और इसका अबद्वतीय हल व्यंजक 𝑋 = 𝐴−1 𝐶 = 𝐵𝐶 द्वारा कदया ½
र्या है
1 −2 −1 3 𝑥 2
⇒ 𝑋 = [0 −1 −1][−2] ⇒ [ ] = [−3]
𝑦 1
2 0 3 5 𝑧 21
इस प्कार ½
𝑥 = 2, 𝑦 = −3, 𝑧 = 21
पेज 9 of 12
Page 23
33A माना 𝑥 = tan𝜃 ⇒ 𝑑𝑥 = 𝑠𝑒𝑐 2 𝜃 𝑑𝜃 ½
𝜋
4 log(1 + 𝑡𝑎𝑛𝜃)
𝐼=∫ . 𝑠𝑒𝑐 2 𝜃𝑑𝜃
0 1 + 𝑡𝑎𝑛2 𝜃
𝜋 𝜋
𝜋
𝐼 = ∫04 𝑙𝑜𝑔 (1 + 𝑡𝑎𝑛𝜃)𝑑𝜃 = ∫04 log[1 + tan( 4 − 𝜃)]𝑑𝜃 1
𝜋
1−tan𝜃
=∫0 log[1 + 1+tan𝜃] 𝑑𝜃
4
1
𝜋
1+tan𝜃+1−tan𝜃
=∫04 log[ 1+tan𝜃 ] d𝜃
𝜋
2
=∫04 log[1+tan𝜃] 𝑑𝜃
𝜋 𝜋
1
= ∫0
4 log2 𝑑𝜃 − ∫0 log[1 + 𝑡𝑎𝑛𝜃] 𝑑𝜃
4
𝜋
log2 × 𝑥]04 − 𝐼 1
𝜋
⇒ 2𝐼 = log2
4
𝜋
⇒ 𝐼 = 8 log2 ½
अथिा
(3sin𝜃−2)cos𝜃 (3sin𝜃−2)cos𝜃
अथिा
33B I = ∫ 5−cos2𝜃−4sin𝜃 𝑑𝜃 = ∫ 5−(1−sin2𝜃)−4sin𝜃 𝑑𝜃
माना sin𝜃 = 𝑡 ⟹ cos𝜃𝑑𝜃 = 𝑑𝑡
(3𝑡−2) ½
𝐼 =∫ 5−(1−𝑡 2 )−4𝑡 𝑑𝑡
(3𝑡−2) 3𝑡−2
=∫ 𝑡 2−4𝑡+4 𝑑𝑡 =∫ (𝑡−2)2 𝑑𝑡
1
3𝑡−2
माना
𝐴 𝐵
(𝑡−2)2
= (𝑡−2) +(𝑡−2)2
3𝑡 − 2 = 𝐴(𝑡 − 2) + 𝐵
दोनों पक्षों में t के र्ुणांकों और बस्थर पदों की तुलना करने पर
𝐴 = 3, −2𝐴 + 𝐵 = −2, 𝐵 = 4
(3sin𝜃−2)cos𝜃 3 4
∫ 5−cos2 𝜃−4sin𝜃 𝑑𝜃 =∫ 𝑡−2 𝑑𝑡 + ∫ (𝑡−2)2 𝑑𝑡 ½ +½
4
= 3log|𝑡 − 2| − 𝑡−2 + 𝐶 1+1
4
= 3log|sin𝜃 − 2| − sin𝜃−2 + C
½
𝑑
34A 𝑦 + (𝑥𝑦) = 𝑥(sin𝑥 + 𝑥)
𝑑𝑥
𝑑𝑦 1
⟹ 𝑦 + (𝑥 𝑑𝑥 + 𝑦 ) = 𝑥(sin𝑥 + 𝑥)
𝑑𝑦
⟹ 2𝑦 + 𝑥 𝑑𝑥 =𝑥(sin𝑥 + 𝑥)
𝑑𝑦 2𝑦
⟹ 𝑑𝑥 + 𝑥 = (sin𝑥 + 𝑥)
यह एक रैबखक अिकल समीकरण 𝑑𝑥 + 𝑃𝑦 = 𝑄 के रूप में है
𝑑𝑦
2 1
जहााँ P=𝑥 ,Q=(sin𝑥 + 𝑥)
2
∫ 𝑑𝑥 2log𝑥 log𝑥 2 2 1
I.F =𝑒 𝑥 = 𝑒 =𝑒 =𝑥
हल होर्ा 𝑦. I.F =∫ 𝑄. 𝐼𝐹𝑑𝑥
𝑦𝑥 2 = ∫ (sin𝑥 + 𝑥)𝑥 2 dx
𝑦𝑥 2 =∫ 𝑠𝑖𝑛𝑥. 𝑥 2 𝑑𝑥 + ∫ 𝑥 3 𝑑𝑥 1
𝑥4
⟹ 𝑦𝑥 2 = −𝑥 2 cos𝑥 + 2∫ 𝑥cos𝑥𝑑𝑥 +4 +C
𝑥4
⟹ 𝑦𝑥 2 = −𝑥 2 cos𝑥 +2 (𝑥𝑠𝑖𝑛𝑥 + cos𝑥)+ 4 + C जो कक आिश्यक हल है 1
पेज 10 of 12
Page 24
अथिा
𝑥 𝑥
34B 2𝑦𝑒 𝑦 𝑑𝑥 + (𝑦 − 2𝑥𝑒 𝑦 ) 𝑑𝑦 = 0 1
𝑥 𝑥
𝑥
𝑑𝑥 2𝑥𝑒 𝑦 −𝑦 2 𝑒 𝑦 −1
𝑦
⟹ 𝑑𝑦 = 𝑥 = 𝑥
2𝑦𝑒 𝑦 2𝑒 𝑦
यह समघातीय अिकल समीकरण है
माना 𝑥 = 𝑣𝑦 =>⟹ 𝑑𝑦 = 𝑣 +𝑦 𝑑𝑦
𝑑𝑥 𝑑𝑣 1
𝑑𝑣 2𝑣𝑒 𝑣 −1
𝑣 +𝑦 𝑑𝑦 = 2𝑒 𝑣
𝑑𝑣 2𝑣𝑒 𝑣 −1 2𝑣𝑒 𝑣 −1−2𝑣𝑒 𝑣
⟹𝑦
𝑑𝑦
= 2𝑒 𝑣 –𝑣 =
2𝑒 𝑣
𝑑𝑣 −1
⟹ 𝑦 𝑑𝑦 =2𝑒 𝑣
1
𝑑𝑦
⟹ 2𝑒 𝑣 𝑑𝑣 = −
𝑦
𝑣 𝑑𝑦
∫ 2𝑒 𝑑𝑣 =−∫ 𝑦
𝑣
⟹ 2𝑒 = −log|𝑦| + C
𝑥
1
⟹ 2𝑒 𝑦 + log|𝑦| = C
जि x=0,y=1, C = 2
𝑥
आिश्यक हल है 2𝑒 𝑦 +log|𝑦| = 2 1
⇒ इस पर कोई बिन्दु है (3 𝜆+1, −𝜆, −1)
𝑥−1 𝑦−0 𝑧+1
35 माना 3
= −1 = 0 =𝜆 ½
बिन्दु के बलए जहााँ 𝑦 = 1 ⇒ 𝜆 = −1
1
⇒ बिन्दु है (−2,1, −1) ½
दो रे खाओं के कदक् हैं ⃗⃗⃗ = 3𝑖̂ − 𝑗̂
𝑚 1
तथा 𝑛⃗⃗ = −2𝑖̂ + 2𝑗̂ + 𝑘̂
½
𝑖̂ 𝑗̂ 𝑘̂
⃗⃗⃗ × 𝑛⃗⃗ = | 3 −1 0| = −𝑖̂ − 3𝑗̂ + 4𝑘̂
𝑚
1
−2 2 1
ज्ञात की जाने िाली रे खा की समीकरण है
𝑟⃗ = (−2𝑖̂ + 𝑗̂ − 𝑘̂ ) + 𝜇(−𝑖̂ − 3𝑗̂ + 4𝑘̂ ) ½
िैकबल्पक हल :
⇒ इस पर कोई बिन्दु है (3𝜆 + 1, −𝜆, −1)
𝑥−1 𝑦−0 𝑧+1
माना 3
= −1 = 0 =𝜆 ½
उस बिन्दु के बलए जहााँ 𝑦 = 1 ⇒ 𝜆 = −1
1
⇒ बिन्दु होर्ा (−2,1, −1)
½
माना कक a, b, c अज्ञात समीकरण के कदक् अनुपात है
ति 3𝑎 − 𝑏 = 0
तथा −2𝑎 + 2𝑏 + 𝑐 = 0 1
𝑎 −𝑏 𝑐 𝑎 𝑏 𝑐 1
हल करने पर = = ⇒ = =
−1 3 4 −1 −3 4
अज्ञात रे खा है :
𝑥+2 𝑦−1 𝑧+1
−1
= −3 = 4 = 𝜇 ½
½
सकदश के रूप में 𝑟⃗ = (−2𝑖̂ + 𝑗̂ − 𝑘̂ ) + 𝜇(−𝑖̂ − 3𝑗̂ + 4𝑘̂ )
पेज 11 of 12
Page 25
खंड -ड़
(3 के स-स्िडी/र्द्ांश-आर्ाररत प्श्न प्त्क्यक
े 4 अंक का )
36 I. यातायात प्िाह स्ितुल्य नहीं है क्योंकक (𝐴, 𝐴) ∉ 𝑅 (या कोई भी प्मुख स्थान स्ियं से 1
जुड़ा नहीं है)
1
II. यातायात प्िाह संक्रामक नहीं है क्योंकक (𝐴, 𝐵) ∈ 𝑅और(𝐵, 𝐸) ∈ 𝑅, परं तु(𝐴, 𝐸) ∉ 𝑅
III. 𝐴. 𝑅 = {(𝐴, 𝐵), (𝐴, 𝐶), (𝐴, 𝐷), (𝐵, 𝐶), (𝐵, 𝐸), (𝐶, 𝐸), (𝐷, 𝐸), (𝐷, 𝐶)}
1
प्ां त = {𝐴, 𝐵, 𝐶, 𝐷}
½ +½
पररसर = {𝐵, 𝐶, 𝐷, 𝐸}
अथिा
III. B.नहीं, ट्रैक़िक प्िाह ककसी फलन को बनरूबपत नहीं करता क्योंकक A की तीन छबियााँ 1+1
हैं ।
37 I. 𝑃(𝑥) = 𝑅(𝑥) − 𝐶(𝑥) = −0.3𝑥 2 + 20𝑥 − (0.5𝑥 2 − 10𝑥 + 150)
=−0.8𝑥 2 + 30𝑥 − 150 1
II. क्रांबतक बिंदओं
ु के बलए 𝑃′ (𝑥) = 0 ⇒ −1.6𝑥 + 30 = 0
30 300
⇒𝑥= = = 18.75
1.6 16 1
III A. अि 𝑃′′ (𝑥) = −1.6
बिशेि रूप से 𝑃′′ (18.75) = −1.6 < 0 1
इसबलए ,क्रांबतक मान 𝑥 = 18.75 अबर्कतम लाभ के अनुरूप है 1
अथिा
III B. क्योंकक 𝑥 िल्िों की संख्या को दशागता है इसबलए बिशेि रूप से 18 िल्ि अबर्कतम
लाभ के अनुरूप है
अबर्कतम लाभ है 𝑃(18) = −0.8 × 182 + 30 × 18 − 150 1
=−259.2 + 540 − 150
=540 − 409.2 = ₹130.80 1
38 घिनाओं को इस प्कार मानें E1: छात्र पहले समूह में है (स्क्रीन पर बिताया र्या समय 4 घंिे से
अबर्क है)
E2: छात्र दूसरे समूह में है (स्क्रीन पर बिताया र्या समय 2 से 4 घंिे है)
E3: छात्र तीसरे समूह में है (स्क्रीन पर बिताया र्या समय 2 घंिे से कम है)
A: छात्र द्वारा बचंता और कम अिर्ारण के लक्षण दशागने की घिना
60 30 10
P(E1) = 100 P(E2) = 100 and P(E3) = 100
80 70 30
P(A|E1) = 100 P(A|E2) = 100 and P(A|E3) = 100
I. P(A)= P(E1) x P(A|E1) + P(E2) x P(A|E2) + P(E3) x P(A|E3)
2
60 80 30 70 10 30 72
= 100 x 100 + 100 x 100 + 100 x 100 = 100 = 72%
𝑃(𝐸1 ∩𝐴)
II. P(E1|A) = 𝑃(𝐴)
60 80
( 𝑥 ) 48 2 2
= 10072100 = 72 = 3
( )
100
पेज 12 of 12
Page 26
Study Materials
Notes
Model Papers Class 6 Notes
Sample Papers Class 7 Notes
Half Yearly Sample Papers Class 8 Notes
Class 9 Notes
Important Resources
Class 10 Notes
Periodic Table
Class 11 Notes
Writing Skills / Formats
Maps of India / World Class 12 Notes
Books and Solutions
NCERT Books
NCERT Book Solutions
HC Verma Chapter Wise Solutions
RD Sharma Solutions
CGBSE Solutions